# Machine Learning Unveils the power law of Finite-Volume Energy Shifts

### v1

@@ Abstract
Here, $n$ is an unknown parameter and needed to be fixed for various samples. The parameter $\mu$ in Eq. (4) is set as $10,~{}20,~{}\dots,120,~{}400,~{}600,~{}800~{}\mathrm{MeV}$ corresponding the range of force $19.73,~{}9.87,\dots,~{}1.64,~{}0.49,~{}0.33,~{}0.25~{}\mathrm{fm}$. The first two ranges of force are compatible with the box size. As analyzed in Ref. [50], the power of $L$ in Eq. (11) decreases when the force range decreases. Moreover, when $\mu$ goes to infinity, i.e. the short-range interaction, the power $n$ comes back to the formula $E_{L}=C_{1}+C_{2}\exp(-C_{3}L)L^{-1}$ deduced from the Lüscher formula [10]. It indicates that no matter the form of the potential, the finite-volume extrapolation formula $E_{L}=C_{1}+C_{2}\exp(-C_{3}L)L^{-1}$ can be reproduced, once the range of force is short enough. On the other hand, when $\mu$ goes to zero, the power goes to infinity, which presents the behavior for the infinity long-range interaction [50]. The range of the force for the first two points of Fig.2 in Ref. [50] are compatible with the box size and cannot reflect the correct finite-volume behavior.
@@
Summary and outlook   Finite-volume extrapolation is an important step for extracting physical observables from lattice calculations. However, it is a significant challenge for the system with long-range interactions, especially when the range of the force is comparable with the lattice box size. Several theoretical works have been put forward to extract the finite-volume extrapolation formula based on the modified Lüscher formula numerically. To obtain an exact extrapolation formula, we employ symbolic regression for both short-range and long-range interactions. The samples are generated by the HLEFT and are fed to symbolic regression. As a benchmark, the finite-volume extrapolation formula for short-range interaction has been successfully regressed, i.e. the exponential form $\exp{(-\kappa L)}$ with the $L^{-1}$ factor in front of it, which is in a good agreement with the theoretical result. Furthermore, we turn to the case of long-range interactions. The regressed results still keep the exponential form, but with the power of $L$ dependent on the range of the force, i.e. the power of $L$ decreases with the decreasing range of the force. When the range of the force becomes sufficiently small, the power converges to $-1$, recovering the short-range formula as expected. This agreement further validates the correctness of our extrapolation formula for long-range forces. Our work represents a significant advancement in leveraging machine learning to probe uncharted territories within particle physics.
@@
Acknowledgments:   We thank Serdar Elhatisari, Feng-Kun Guo, Ulf-G. Meißner, Andreas Nogga, Akaki G. Rusetsky for their useful discussions on the finite volume dependence from the theoretical side. This work is supported in partly by the NSFC under Grants Nos. 12375072, 12375073, 12275259, and 12135011. The work of Q.W. is also supported by Guangdong Provincial funding under Grant No. 2019QN01X172. B.N. Lu is also supported by NSAF No.U2330401.

### v2

@@ Abstract
Here, $n$ is an unknown parameter and needs to be fixed for various samples. The parameter $\mu$ in Eq. (4) is set as $10,~20,~\dots,120,~400,~600,~800~\mathrm{MeV}$ corresponding to the range of force $19.73,~9.87,\dots,~1.64,~0.49,~0.33,~0.25~\mathrm{fm}$. The first two ranges of force are compatible with the box size. As analyzed in the SM, the power of $L$ in Eq. (11) decreases when the force range decreases. Moreover, when $\mu$ goes to infinity, i.e., the short-range interaction, the power $n$ comes back to the formula $E_{L}=C_{1}+C_{2}\exp(-C_{3}L)L^{-1}$ deduced from the Lüscher formula. [11] It indicates that no matter the form of the potential, the finite-volume extrapolation formula $E_{L}=C_{1}+C_{2}\exp(-C_{3}L)L^{-1}$ can be reproduced once the range of force is short enough. On the other hand, when $\mu$ goes to zero, the power goes to infinity, which presents the behavior for the infinite long-range interaction (see the SM). The range of the force for the first two points of Fig.S2 of SM is compatible with the box size and cannot reflect the correct finite-volume behavior.
@@
5. Summary and Outlook.   Finite-volume extrapolation is an important step for extracting physical observables from lattice calculations. However, it is a significant challenge for systems with long-range interactions, especially when the range of the force is comparable to the lattice box size. Several theoretical works have been put forward to extract the finite-volume extrapolation formula based on the modified Lüscher formula numerically. To obtain an exact extrapolation formula, we employ symbolic regression for both short-range and long-range interactions. The samples are generated by the HLEFT and are fed to symbolic regression. As a benchmark, the finite-volume extrapolation formula for short-range interaction has been successfully regressed, i.e., the exponential form $\exp{(-\kappa L)}$ with the $L^{-1}$ factor in front of it, which is in good agreement with the theoretical result. Furthermore, we turn to the case of long-range interactions. The regressed results still keep the exponential form, but with the power of $L$ dependent on the range of the force, i.e., the power of $L$ decreases with the decreasing range of the force. When the range of the force becomes sufficiently small, the power converges to 1-1- 1, recovering the short-range formula as expected. This agreement further validates the correctness of our extrapolation formula for long-range forces. Our work represents a significant advancement in leveraging machine learning to probe uncharted territories within particle physics.
@@
Acknowledgments.   We thank Serdar Elhatisari, Feng-Kun Guo, Ulf-G. Meißner, Andreas Nogga, and Akaki G. Rusetsky for their useful discussions on the finite volume dependence from the theoretical side. This work was supported in part by the National Natural Science Foundation of China (Grant Nos. 12375072, 12375073, 12275259, and 12135011). The work of Q.W. was also supported by Guangdong Provincial Funding (Grant No. 2019QN01X172). B.N. Lu is also supported by the National Security Academic Fund (Grant No.U2330401).